# Long-Duration Station-Keeping Strategy for Cislunar Spacecraft Formations

### v1

@@ 4.4 Maneuver Time Placement
The performance of the absolute control, relative control, and safety filter are heavily dependent on the placement of maneuver times. The nature of NRHO trajectories dictates that maneuvers be performed near apoapsis crossing, where the dynamics are far less sensitive so as to not amplify navigation and control execution errors [12, 14]. Applying maneuvers within the time intervals corresponding to true anomalies between $160\deg$ and $200\deg$, a duration of 3.6$\sim$ 3.6∼ 3.6 days, provides ample maneuver freedom without incurring large errors due to dynamical sensitivity. This duration is also sufficiently long to allow for the resolution of and recovery from MMEs by ground operators for absolute control maneuvers, which have the greatest magnitude and therefore the most severe effect when missed. It is also best for absolute maneuvers, which are applied once per orbit, to be performed as early as possible in this interval, so that the ensuing relative maneuvers can be used to clean up formation dispersions introduced by the absolute maneuvers. This ensures that the magnitude of the relative dispersions at perilune crossing, the most sensitive and chaotic region, is low. Thus, absolute maneuvers are performed at times corresponding to a true anomaly of $160\deg$ and several relative maneuvers are performed up to and including a true anomaly of $200\deg$. The number of relative maneuvers is a design parameter dependent on desired $\Delta V$ expenditure, which increases with maneuver frequency, and tracking performance, which improves with maneuver frequency.
@@ 4.5 Virtual Chief Placement
The choice of absolute state to represent the formation can have a dramatic effect on the $\Delta V$ consumption of the formation and the extent to which this consumption is spread among the agents. For example, a virtual chief can be selected, for which the deputy spacecraft maintain a relative orbit with respect to, such that the overall fuel expenditure or distribution of fuel expenditures is minimized. This has been demonstrated in circular orbits [41] and eccentric orbits [42] for Keplerian motion. Especially for large swarms, it may be detrimental to have a specific spacecraft operate as the chief, since this relocates the disturbances on the chief to the deputies, removing relative station-keeping burden from the chief and placing it on the deputies. In general, the $\Delta V$ associated with the choice of virtual chief is difficult to determine since both absolute and relative $\Delta V$ expenditure is affected. However, a simple and effective strategy that distributes relative station-keeping burden among the formation is to place the virtual chief at the mean of the state estimates. If the collection of absolute state estimates from the navigation system are given by $\{\hat{\bm{x}}_{k}\}_{k=1}^{M}$, then the virtual chief state and deputy states are determined by
@@ 5 Navigation Simulation
To accurately represent and simulate the performance of the station-keeping algorithms, realistic navigation algorithms are implemented. Navigation updates are given in the form of absolute and relative pseudo-range and pseudo-range rates, representing Deep Space Network (DSN) and intersatellite (IS) measurements and fed to an Extended Kalman Filter (EKF). The measurement architectures explored include a cluster architecture, whereby measurements include the absolute range of a hub spacecraft and the relative ranges between all spacecraft, a fully connected architecture, whereby measurements include the absolute ranges of all spacecraft and relative ranges between all spacecraft, and a hub-spoke architecture, whereby measurements include absolute range of a hub spacecraft and relative range between the hub and spoke spacecraft. Absolute ranging is scheduled to occur at -72 hours, -48 hours, -7 hours and +12 hours with respect to an absolute maneuver time, with 10 measurements taken over the span of one hour, in line with the measurement frequency proposed in Ref. 14. In addition, relative ranging is scheduled for the same time as absolute ranging and 1 hour before any relative maneuver time. Relative navigation can be performed autonomously, provided the availability of STMs onboard the spacecraft.
Though the navigation performance is dependent on the architecture and number of agents, typically absolute and relative errors are on the order of 1 km and 10 mm/s and 10 m and .01 mm/s respectively preceding maneuvers. Tab. 1 contains the relevant navigation parameters applied to all ensuing test cases.
@@ 6 Test Scenarios and Results
    A two spacecraft fully connected formation with 10$\sim$ 10∼ 10 km separation under a noisy configuration.
@@
    A five spacecraft cluster formation with 1$\sim$ 1∼ 1 km relative separation under a quiet configuration.
@@
    A nine spacecraft hub-spoke formation with 100$\sim$ 100∼ 100 km relative separation under a noisy configuration.
@@
Fixed Magnitude Execution Error .01 [mm/s] .3 [mm/s]
Proportional Magnitude Execution Error .1 [%] 1 [%]
Fixed Pointing Execution Error .01 [mm/s] .3 [mm/s]
Proportional Magnitude Execution Error .001 [rad] .01 [rad]
@@
Minimum Safe Separation 100 [m] 100 [m]
@@
    Randomize the reflective coefficient associated with each agent in accordance with the SRP errors in Tab. 2. Propagate the agents until the next maneuver time. Propagate the navigation filter with measurements applied at relevant times.
@@
    Remove maneuvers that violate the minimum impulse bit defined in Tab. 2. Apply maneuver execution errors from Gate’s model [38] with errors from Tab. 2. If a random number is below the MME rate, then randomly sample a propulsion cutoff time. All maneuvers for the given agent are set to zero for the remainder of this cutoff time. Add the resulting $\Delta V$ to the state of each agent.
Figs. 10, 12, and 15 show the performance of the proposed strategy for each respective test case, including navigation, relative and absolute tracking, $\Delta V$ histories, and trajectories in the VNB frame and local toroidal coordinates. Several performance metrics are summarized in Tab. 3, including yearly $\Delta V$ expenditure, separation ranges, and computation times [^1] Experiments were performed on an Intel i7‑10750H CPU. The code was implemented and executed in MATLAB R2024a. The algorithms made use of MOSEK [43] to optimize and YALMIP to model [44]. [^2] Similar absolute control strategy implementations (e.g., Ref. 14) have achieved lower computation times.. Relative separation ranges denote the absolute minimum and maximum separation between the agents in the formation over all trials. The activation rate of the CBF-SIP represents the percentage of time that at least one iteration of the CBF optimization routine occurs, i.e. the desired $\Delta V$ is not already passively-safe.
    Metric Test Case 1, Fig. 10 Test Case 2, Fig. 12 Test Case 3, Fig. 15
Monte Carlo Trials 20 20 20
Computed $\Delta V_{abs}$, [cm/s/year] 12.33±3.4512.33$\pm$ 3.4512.33 ± 3.45 10.41±3.2710.41$\pm$ 3.2710.41 ± 3.27 11.37±3.9311.37$\pm$ 3.9311.37 ± 3.93
Total Executed $\Delta V$ per agent, [cm/s/year] $\left[\begin{array}[]{@{}l@{}}21.06\pm 3.75\\ 21.30\pm 3.38\end{array}\right]$ $\left[\begin{array}[]{@{}l@{}}13.15\pm 3.74,13.60\pm 3.73\\ 14.62\pm 4.21,13.17\pm 3.46\\ 14.56\pm 3.76\end{array}\right]$ $\left[\begin{array}[]{@{}l@{}}25.60\pm 4.73,36.89\pm 4.21\\ 43.91\pm 6.28,34.93\pm 4.54\\ 42.64\pm 6.20,45.61\pm 7.45\\ 50.22\pm 8.19,45.99\pm 5.95\\ 48.43\pm 4.71\end{array}\right]$
Relative Separation Ranges [km] [0.77,41.19] [0.098,7.49] [3.24,518]
CBF-SIP Activation Rate [%] 0.1 16.43 0
Absolute Maneuver Computation Time [sec] 35.63±13.6235.63$\pm$ 13.6235.63 ± 13.62 38.59±15.4838.59$\pm$ 15.4838.59 ± 15.48 38.91±16.3238.91$\pm$ 16.3238.91 ± 16.32
Relative Maneuver Computation Time [millisec] 0.46±0.290.46$\pm$ 0.290.46 ± 0.29 0.50±0.260.50$\pm$ 0.260.50 ± 0.26 1.9±161.9$\pm$ 161.9 ± 16
CBF-SIP Computation Time [sec] 0.02±0.460.02$\pm$ 0.460.02 ± 0.46 0.20±0.260.20$\pm$ 0.260.20 ± 0.26 0.40±0.190.40$\pm$ 0.190.40 ± 0.19
@@ 7 Discussion
For all trials, the strategy demonstrates successful station-keeping for the formations with low to moderate $\Delta V$ expenditure, adequate reference tracking, low computational cost, and guarantees of passive safety. Over the cumulative 53 years of simulated absolute and relative station-keeping, the absolute maneuver and CBF-SIP schemes never fail to converge and absolute and relative bounded motion is always achieved. Moreover, no collision risk is detected, with the absolute lowest separation occuring for Test Case 2 being 980 meters.
@@ 7.1 Absolute Control Performance
Absolute reference tracking is achieved on the level of hundreds of kilometers and absolute station-keeping costs are on the order of tens of centimeters per second. These figures match the findings of other studies of absolute NRHO control [12, 14]. The two spacecraft and five spacecraft formation test cases demonstrated the lowest computed absolute $\Delta V$. This is likely attributable to the lower absolute navigation errors achieved as a result of the fully connected and cluster measurement architecture whereby the two spacecraft both receive absolute ranging measurements. This demonstrates an important benefit associated with cislunar formation flight, whereby multiple sources of absolute measurements, when combined with relative measurements, contribute to improved absolute state knowledge. Previous studies, such as Refs. 12, 14, have found that absolute state knowledge error is the primary driver of $\Delta V$ consumption and absolute reference tracking, suggesting that distributed navigation architectures over cislunar formations can provide benefits to absolute control of NRHOs. A decrease in computed absolute maneuver expenditure could be additionally attributable to relative station-keeping, which cleans up after missed absolute maneuvers.
@@ 7.2 Relative Control Performance
The relative control concept, which leverages STMs generated about the reference trajectory for dynamics representation, QPROs as relative references, and a simple control law, demonstrates fine level of relative maneuvering under high-fidelity cislunar simulation. These features make the relative control law highly desirable in comparison to computing absolute maneuvers for each agent, which is incapable of providing precise relative positioning.
The formation experiences the largest relative dispersions at perilune, the region corresponding to the greatest relative separations and most chaotic dynamics. Even under low levels of noise, it is difficult to ensure good tracking at perilune as a result of only performing maneuvers at apolune. This can be compared with the operations of the PROBA-3 spacecraft [45], which operates in a highly eccentric orbit about the Earth and performs high accuracy relative control only at apoapsis. The greatest contributor to this relative dispersion under normal conditions is the dynamical infeasibility of the QPRO. Since the QPRO is generated about a reference trajectory, and absolute control induces hundreds of kilometers of dispersion from this reference, the linearized dynamics from which the QPRO is generated degrade in accuracy. This fact has not been appropriately considered in previous studies of cislunar formation flight, which typically use simplified models and assume perfect absolute reference knowledge and tracking.
@@
Relative maneuver expenditure is also strongly correlated with the size of the formation, with Test Case 3, with hundreds of kilometers of separation, incurring the greatest $\Delta V$ expenditure. This can be attributed to the fact that the linearized dynamics, which are used by the relative control law, become less accurate at higher separations.
The proposed control law, though simple and efficient for station-keeping tasks, is likely inadequate for more complex maneuvering tasks, such as phase changing, approach trajectory generation, and QPRO reconfiguration. However, in tandem with more advanced relative control techniques, such as those proposed in Refs. 46, 20, advanced maneuvering strategies could be achieved during operation.
@@ 7.3 Safety Filter Performance
A key feature of the proposed CBF-SIP is that computation only occurs if the formation enters an unsafe condition, leveraging the rarity of near-miss collision events. As a result, both Test Case 1 and Test Case 3 feature no activations of the CBF-SIP, meaning runtime is only spent verifying that ensuing trajectories are safe. In contrast, Test Case 2 features 20 activations of the CBF-SIP, given the tighter formation.
The ability of the CBF-SIP to guarantee safety is demonstrated in Fig. 17, which originates from Test Case 2. When the CBF-SIP is not used, a near-miss event occurs with the spacecraft trajectories coming to within 38 meters of each other following a missed station-keeping maneuver. With the CBF-SIP active, the near-miss event is avoided and the closest approach occurs at 100 meters, the prescribed bound.
@@ 8 Conclusion
This work presents a guidance and control strategy for long duration station-keeping in cislunar near-rectilinear halo orbits and demonstrates safe and efficient deployment in high-fidelity scenarios. The strategy represents one of the first demonstrations of a complete GN&C architecture applied to cislunar formation flight in NRHOs under high-fidelity simulation. The combination of state-of-the-art absolute reference tracking strategies with a simple control law for QPRO tracking is achieved through the introduction of a control barrier function safety filter, which allows for decoupling of the absolute and relative control formulations. Relative navigation, relative control, and the CBF-SIP, are computationally efficient and rely solely on linear relative dynamical knowledge, meaning that these aspects of formation-keeping could be performed autonomously, provided ample storage for STM data. In this respect, the proposed framework would introduce minor amounts of additional operational burden compared to a single spacecraft mission.
Finally, it is emphasized that though this work has focused on the relevant application of the 9:2 $L_{2}$ NRHO, the proposed strategy is generalizable to different cislunar regimes. Future work includes the systematic verification of the GN&C strategy to other orbits of interest, including distant retrograde orbits, butterfly orbits, Lyapunov orbits, and other $L_{1}/L_{2}$ Halo orbits, the inclusion of advanced reconfiguration strategies, such as those proposed in Ref. 21, and further analysis of strategies for generating longer-duration QPROs.

### v2

@@ 4.4 Maneuver Time Placement
The performance of the absolute control, relative control, and safety filter are heavily dependent on the placement of maneuver times. The nature of NRHO trajectories dictates that maneuvers be performed near apoapsis crossing, where the dynamics are far less sensitive so as to not amplify navigation and control execution errors [davis2022orbit, shimane2025output]. Applying maneuvers within the time intervals corresponding to true anomalies between $160\deg$ and $200\deg$, a duration of $\sim 3.6$ days, provides ample maneuver freedom without incurring large errors due to dynamical sensitivity. This duration is also sufficiently long to allow for the resolution of and recovery from MMEs by ground operators for absolute control maneuvers, which have the greatest magnitude and therefore the most severe effect when missed. It is also best for absolute maneuvers, which are applied once per orbit, to be performed as early as possible in this interval, so that the ensuing relative maneuvers can be used to clean up formation dispersions introduced by the absolute maneuvers. This ensures that the magnitude of the relative dispersions at perilune crossing, the most sensitive and chaotic region, is low. Thus, absolute maneuvers are performed at times corresponding to a true anomaly of $160\deg$ and several relative maneuvers are performed up to and including a true anomaly of $200\deg$. The number of relative maneuvers is a design parameter dependent on desired $\Delta V$ expenditure, which increases with maneuver frequency, and tracking performance, which improves with maneuver frequency.
@@ 4.5 Virtual Chief Placement
The choice of absolute state to represent the formation can have a dramatic effect on the $\Delta V$ consumption of the formation and the extent to which this consumption is spread among the agents. For example, a virtual chief can be selected, for which the deputy spacecraft maintain a relative orbit with respect to, such that the overall fuel expenditure or distribution of fuel expenditures is minimized. This has been demonstrated in circular orbits [LIPPE2021162] and eccentric orbits [lippevcplacement] for Keplerian motion. Especially for large swarms, it may be detrimental to have a specific spacecraft operate as the chief, since this relocates the disturbances on the chief to the deputies, removing relative station-keeping burden from the chief and placing it on the deputies. In general, the $\Delta V$ associated with the choice of virtual chief is difficult to determine since both absolute and relative $\Delta V$ expenditure is affected. However, a simple and effective strategy that distributes relative station-keeping burden among the formation is to place the virtual chief at the mean of the state estimates. If the collection of absolute state estimates from the navigation system are given by $\{\hat{\bm{x}}_{k}\}_{k=1}^{M}$, then the virtual chief state and deputy states are determined by
@@ 5 Navigation Simulation
To accurately represent and simulate the performance of the station-keeping algorithms, realistic navigation algorithms are implemented. Navigation updates are given in the form of absolute and relative pseudo-range and pseudo-range rates, representing Deep Space Network (DSN) and intersatellite (IS) measurements and fed to an Extended Kalman Filter (EKF). The measurement architectures explored include a cluster architecture, whereby measurements include the absolute range of a hub spacecraft and the relative ranges between all spacecraft, a fully connected architecture, whereby measurements include the absolute ranges of all spacecraft and relative ranges between all spacecraft, and a hub-spoke architecture, whereby measurements include absolute range of a hub spacecraft and relative range between the hub and spoke spacecraft. Absolute ranging is scheduled to occur at -72 hours, -48 hours, -7 hours and +12 hours with respect to an absolute maneuver time (true anomaly of $160\deg$), with 10 measurements taken over the span of one hour, in line with the measurement frequency proposed in Ref. shimane2025output. In addition, relative ranging is scheduled for the same time as absolute ranging and 1 hour before any relative maneuver time for the fully connected architecture, every 24 hours for the cluster architecture, and every 6 hours for the hub and spoke architecture. Tab. 1 contains the relevant navigation parameters applied to all ensuing test cases.
@@ 6 Test Scenarios and Results
    A two spacecraft fully connected formation with $\sim 10$ km separation under a noisy configuration.
@@
    A five spacecraft cluster formation with $\sim 1$ km relative separation under a quiet configuration.
@@
    A nine spacecraft hub-spoke formation with $\sim 100$ km relative separation under a noisy configuration.
@@
Fixed Magnitude Execution Error 0.01 [mm/s] 0.3 [mm/s]
Proportional Magnitude Execution Error 0.1 [%] 1 [%]
Fixed Pointing Execution Error .01 [mm/s] 0.3 [mm/s]
Proportional Pointing Execution Error 0.001 [rad] 0.01 [rad]
@@
Minimum Safe Separation 100 [m] 1000 [m]
@@
    Metric Test Case 1, Fig. 10 Test Case 2, Fig. 12 Test Case 3, Fig. 15
Monte Carlo Trials 10 10 10
Computed $\Delta V_{abs}$, [cm/s/year] $53.99\pm 7.77$ $9.08\pm 2.61$ $52.63\pm 5.44$
Total Executed $\Delta V$ per agent, [cm/s/year] $\left[\begin{array}[]{@{}l@{}}81.39\pm 9.42\\ 75.34\pm 9.49\end{array}\right]$ $\left[\begin{array}[]{@{}l@{}}29.65\pm 7.47,23.63\pm 5.62\\ 24.76\pm 4.33,23.85\pm 5.91\\ 24.46\pm 5.64\end{array}\right]$ $\left[\begin{array}[]{@{}l@{}}75.50\pm 5.30,82.76\pm 7.15\\ 88.03\pm 10.17,89.99\pm 22.93\\ 96.85\pm 17.91,89.43\pm 8.76\\ 93.08\pm 11.58,99.58\pm 10.78\\ 100.29\pm 7.74\end{array}\right]$
Relative Separation Ranges [km] [1.04,57.90] [0.099,7.34] [2.38,523.98]
CBF-SIP Activation Rate [%] 2.8 29.13 0.01
Absolute Maneuver Computation Time [sec] $63.40\pm 29.65$ $45.34\pm 18.89$ $84.64\pm 43.03$
Relative Maneuver Computation Time [millisec] $0.49\pm 0.30$ $0.60\pm 0.77$ $1.44\pm 4.92$
CBF-SIP Computation Time [sec] $0.07\pm 0.32$ $0.89\pm 1.36$ $0.89\pm 0.58$
@@
    Uniformly sample a reflective coefficient associated with each agent in accordance with the SRP errors in Tab. 2. Propagate the agents until the next maneuver time. Propagate the navigation filter with measurements applied at relevant times.
@@
    Remove maneuvers that violate the minimum impulse bit defined in Tab. 2. Apply maneuver execution errors from Gate’s model [Gates1963] with errors from Tab. 2. If a uniformly sampled random number is below the MME rate, then uniformly sample a propulsion cutoff time between the minimum and maximum bounds. All maneuvers for the given agent are set to zero for the remainder of this cutoff time. Add the resulting $\Delta V$ to the state of each agent.
Figs. 10, 12, and 15 show the performance of the proposed strategy for each respective test case, including navigation, relative and absolute tracking, $\Delta V$ histories, and trajectories in the VNB frame and local toroidal coordinates. Several performance metrics are summarized in Tab. 3, including yearly $\Delta V$ expenditure, separation ranges, and computation times [^1] Experiments were performed on an Intel i7‑10750H CPU. The code was implemented and executed in MATLAB R2024a. The algorithms made use of MOSEK [mosek] to optimize and YALMIP [Lofberg2004] to model. [^2] Similar absolute control strategy implementations (e.g., Ref. shimane2025output) have achieved lower computation times.. Relative separation ranges denote the absolute minimum and maximum separation between the agents in the formation over all trials. The activation rate of the CBF-SIP represents the percentage of times that at least one iteration of the CBF optimization routine occurs, i.e. the desired $\Delta V$ is not already passively-safe.
@@ 7 Discussion
For all trials, the strategy demonstrates successful station-keeping for the formations with low to moderate $\Delta V$ expenditure, adequate reference tracking, low computational cost, and guarantees of passive safety. Over the cumulative 25 years of simulated absolute and relative station-keeping, the absolute maneuver and CBF-SIP schemes never fail to converge and absolute and relative bounded motion is always achieved. Moreover, no significant collision risk is detected, with the absolute lowest separation being 99 meters for Test Case 2 and 1040 meters for Test Case 1.
@@ 7.1 Absolute Control Performance
Absolute reference tracking is achieved on the level of hundreds of kilometers and absolute station-keeping costs are on the order of tens of centimeters per second. These figures match the findings of other studies of absolute NRHO control [davis2022orbit, shimane2025output]. Lower computed absolute station-keeping maneuver magnitudes could be attributable to the connected navigation architectures, which likely produce more accurate navigation estimates. This demonstrates an important benefit associated with cislunar formation flight, whereby multiple sources of absolute measurements, when combined with relative measurements, produce improved absolute state knowledge. Previous studies, such as Refs. davis2022orbit, shimane2025output, have found that absolute state knowledge error is the primary driver of $\Delta V$ consumption and absolute reference tracking, suggesting that distributed navigation architectures over cislunar formations can provide benefits to absolute control of NRHOs. A decrease in computed absolute maneuver expenditure could be additionally attributable to offloading to relative station-keeping, which cleans up after missed absolute maneuvers.
@@ 7.2 Relative Control Performance
The relative control concept, which leverages STMs generated about the reference trajectory for dynamics representation, QPROs as relative references, and a simple control law, demonstrates effective relative station-keeping under high-fidelity cislunar simulation. These features make the relative control law highly desirable in comparison to computing absolute maneuvers for each agent, which is incapable of providing precise relative positioning.
The formation experiences the largest relative dispersions at perilune, the region corresponding to the greatest relative separations and most chaotic dynamics. Even under low levels of noise, it is difficult to ensure good tracking at perilune as a result of only performing maneuvers at apolune. This can be compared with the operations of the PROBA-3 spacecraft [Garcia2017_PROBA3_FDS], which operates in a highly eccentric orbit about the Earth and performs high accuracy relative control only at apogee. The greatest contributor to this relative dispersion under normal conditions is the dynamical infeasibility of the QPRO. Since the QPRO is generated about a reference trajectory, and absolute control induces hundreds of kilometers of dispersion from this reference, the linearized dynamics from which the QPRO is generated degrade in accuracy. This fact has not been appropriately considered in previous studies of cislunar formation flight, which typically use simplified models and assume perfect absolute reference knowledge and tracking. Another implication of this finding is the degradation in the validity of the LTCs over the noisy, sensitive, and dispersed environment of the 9:2 NRHO.
@@
Relative maneuver expenditure is also strongly correlated with the size of the formation, as seen with Test Case 3, where hundreds of kilometers of separation incur the greatest $\Delta V$ expenditure. This can be attributed to the fact that the linearized dynamics, which are used by the relative control law, become less accurate at higher separations.
@@ 7.3 Safety Filter Performance
A key feature of the proposed CBF-SIP is that optimization routines are only called if the formation enters an unsafe condition, leveraging the rarity of near-miss collision events. As a result, Test Case 3 features almost no activations of the CBF-SIP, meaning runtime is only spent verifying that ensuing trajectories are safe. In contrast, Test Case 2 features a 29% activation rate of the CBF-SIP, given the tighter formation, and yet retains long-term stability over the duration of flight.
The ability of the CBF-SIP to guarantee safety is demonstrated in Fig. 17, which originates from Test Case 2. When the CBF-SIP is not active, a near-miss event occurs with the spacecraft trajectories coming to within 38 meters of each other following a missed station-keeping maneuver. With the CBF-SIP active, the CBF appropriately corrects the maneuvers such that the near-miss event is avoided and the closest approach occurs at 100 meters, the prescribed bound.
@@ 8 Conclusion
This work presents a guidance and control strategy for long duration station-keeping in cislunar near-rectilinear halo orbits and demonstrates safe and efficient deployment in high-fidelity scenarios. The strategy represents one of the first demonstrations of a complete GN&C architecture applied to cislunar formation flight in NRHOs under high-fidelity simulation. The combination of state-of-the-art absolute reference tracking strategies with a simple control law for QPRO tracking is achieved through the introduction of a control barrier function safety filter, which allows for decoupling of the absolute and relative control formulations. Relative navigation, relative control, and the CBF-SIP, are computationally efficient and introduce minimal additional computational burden, meaning that these aspects of formation-keeping could be performed autonomously. In this respect, the proposed framework has the potential to introduce minor amounts of additional operational burden compared to a single spacecraft mission.
Finally, it is emphasized that though this work has focused on the relevant application of the 9:2 $L_{2}$ NRHO, the proposed strategy is generalizable to other multi-body orbits. Future work includes the systematic verification of the GN&C strategy to other orbits of interest, including distant retrograde orbits, butterfly orbits, Lyapunov orbits, and other $L_{1}/L_{2}$ Halo orbits, the inclusion of advanced reconfiguration strategies, such as those proposed in Refs. takubo2025safeoptimalnspacecraftswarm, khoury_relative_2024, and further analysis of strategies for generating longer-duration QPROs. In addition, the inclusion of greater sources of noise, such as larger solar radiation pressure effects and attitude desaturation events, would be beneficial to include for further validation.